OpenStudy (anonymous):

definite integral. e^x (sinx). upper pi. lower pi/2. show steps.

OpenStudy (anonymous):

use product rule.. do you know the product rule of integration? If not.. let me know

OpenStudy (anonymous):

(fg)' = fg' + gf' ??

OpenStudy (anonymous):

dude that is for differentiation.. your question is to integrate, isn't it? So I assume you don't know.. I would rather type here..

OpenStudy (anonymous):

i dunno it.

OpenStudy (anonymous):

can reduction formula be used?

OpenStudy (anonymous):

use by parts method or integration table

OpenStudy (dumbcow):

product rule he is referring to is "integration by parts" i believe

OpenStudy (anonymous):

ok, IBP. i know that.

OpenStudy (dumbcow):

you'll have to do it twice, be careful with the signs

OpenStudy (anonymous):

alien when you integrate you'll get something like this \[\int sin(x)e^x=f(x)\pm g(x)..... -\int sin(x)e^x\]

OpenStudy (anonymous):

then all you do is add the integral to the otherside and divide by two

OpenStudy (anonymous):

so i got [ e^pi - e^(pi/2) ] / 2

OpenStudy (anonymous):

...for the final.

OpenStudy (dumbcow):

yep, thats correct you can always check your answer with wolfram too http://www.wolframalpha.com/input/?i=intgrate+e^x+*sin%28x%29+from+pi%2F2+to+pi

OpenStudy (anonymous):

hmm, do u guys think gold, silver, bronze medal system would be a good suggestion? and have something like 3 bronze = 1 silver and 3 silver = 1 gold?

OpenStudy (dumbcow):

maybe..it might give more feedback but could be a hassle to implement

OpenStudy (anonymous):

ya it would require their programmers to do more work :) ... also need more storage for who has what